OpenStudy (anonymous):

A Preschool class has 3-year olds and 4-year olds. There are 17 children in the class, and the total of their ages is 63. Which system of linear equations represents the problem

OpenStudy (anonymous):

(3x) + (4y) = 63 x+ y = 17

OpenStudy (anonymous):

By substitution... (3*(17-y))+(4y) = 63 51-3y+4y=63 y=63-51 y=12

OpenStudy (anonymous):

Since x+y=17, for y=12 we have x=5. X represents the number of 3 year olds and y represents 4 year olds